OpenStudy (anonymous):

Consider the logarithm function \(log=exp^{-1}\) on \(\mathbb{R}\). Prove: d) \(log : {\mathbb{R}}_{+} \rightarrow \mathbb{R} \) is strictly monotonically increasing, and thus injective.

OpenStudy (experimentx):

\[ f(x) = \ln x\\ f'(x) = {1 \over x} \\ f'(R_+) = +ve \text{ hence it is increasing} \]

OpenStudy (experimentx):

for any two +ve number 'a' and 'b' such that a = b exp^-1(a) = exp^-1(b) => log(a) = log(b) and for any two real numbers log(a) = log(b) => e^log(a) = e^log(b) => a = b since it is strictly increasing function (sinve the slope > 0), we can say that it is injective function

OpenStudy (anonymous):

thanks experimentX \[ln x\] is the reverse of \[e^{x}\] right ?

OpenStudy (experimentx):

inverse

OpenStudy (anonymous):

aah ok sorry my english not so good :) thank you

OpenStudy (experimentx):

yw